# A finite element solver for a thermodynamically consistent electrolyte model

### v1

@@ 3.2 Variational formulation and boundary implementation
$V^{(w)}$ =$=\left\{u\in H^{1}(\Omega)\ \middle|\ u=g_{w}\text{ on }\Gamma_{D}^{(w)}\right\},\quad\quad V_{0}={\left\{u\in H^{1}(\Omega)\ \middle|\ u=0\text{ on }\partial\Omega\right\}},$ (37)
@@ 4 Convergence study and validation
$T$ $=293.75\,\text{K},\quad n^{\text{ref}}=55\cdot N_{A}\cdot 10^{-3}\,m^{-3},\quad p^{\text{ref}}=1.01325\cdot 10^{5}\,\text{Pa},$
$L^{\text{ref}}$ $=20\cdot 10^{-9}\,\text{m},\quad\chi=80,\quad\lambda^{2}=8.553\cdot 10^{-6},\quad a^{2}=7.5412\cdot 10^{-4}.$
@@ 4.1 Convergence study
Figure 2 illustrates the error in the $e_{2}$- and $e_{\infty}$-norms as a function of the number of elements ($n_{x}$) on a logarithmic scale. The $e_{2}$-norm exhibits third-order convergence, while the $e_{\infty}$-norm demonstrates second-order convergence. The convergence of both error norms toward zero confirms the solver’s consistency and its ability to approximate the reference solution across the entire domain with increasing resolution.
@@ 5.1.1 Comparison to the classical NP model
The thermodynamically consistent electrolyte model utilized in this paper (hereafter referred to as the DGM model), following [9], addresses key limitations of the classical Nernst–Planck (NP) model and its Poisson–Boltzmann formulation. One central distinction is the proper coupling of diffusion fluxes. In the DGM model, only $N-1$ fluxes are treated as independent, while the flux of the remaining species is obtained from the constraint
@@
which ensures mass conservation and consistency with the barycentric frame. Consequently, the model formulation leads to $N-1$ coupled diffusion equations instead of $N$ uncoupled ones, as seen in the classical NP framework.
@@
Figure 4 compares the anion atomic fraction profiles computed using the classical NP and thermodynamically consistent DGM models under varying applied potential differences. It clearly illustrates that the NP model yields unrealistic concentrations that violate physical bounds, especially near interfaces. In contrast, the DGM model maintains the solution within admissible limits, preventing divergence and ensuring consistency with thermodynamic constraints. Additionally, Figure 5 presents a broader comparison of electric potential ($\varphi$), pressure ($p$), and atomic fraction ($y_{\alpha}$) profiles at applied voltages $\delta\varphi=1$ and $\delta\varphi=10$. As the applied voltage increases, the classical NP model produces diverging ion concentrations, especially for anions and the neutral solvent, while the DGM model exhibits saturation behavior, limiting atomic fractions to the physical range. These discrepancies become more pronounced under strong electric fields, where the DGM model better reflects the physical blocking of ionic species.
@@ 5.2.3 Effect of bulk modulus
Figure 12 illustrates the effect of the bulk modulus $K$—which controls the compressibility of the electrolyte—on the one-dimensional electrostatic charge and differential double layer capacitance. The simulations are carried out for increasing compressibility, corresponding to decreasing $K$ values: $K=5\times 10^{4},4\times 10^{4},3\times 10^{4}$, $2\times 10^{4}$, with the incompressible limit ($K\to\infty$). All cases assume a fixed low molarity of $M=0.01\,\text{mol}/\text{L}$. Fig. 12(a) shows that the total electrostatic charge $Q$ increases monotonically with the applied potential difference $\delta\varphi$ for all values of $K$. As expected, compressibility affects the system’s ability to accumulate space charge: lower values of $K$ (i.e., higher compressibility) lead to slightly higher stored charge at the same potential. The differences are small near $\delta\varphi=0$, but they grow with increasing magnitude of the applied potential. As $K$ increases, the results converge smoothly toward the incompressible solution, confirming the theoretical expectation. The more pronounced effect of compressibility appears in the capacitance behavior, shown in Fig. 12(b). For the incompressible case and high $K$ values, the capacitance exhibits a distinct camel-shaped profile with two symmetric peaks and a local minimum at $\delta\varphi=0$. This shape reflects the steric saturation of ions near the electrodes at high field strengths. However, as compressibility increases (i.e., with lower $K$), the profile evolves: the twin humps become broader and less pronounced, and for sufficiently low $K$ (e.g., $K=2\times 10^{4}$), the capacitance begins to increase again beyond the classical humps, deviating from the camel shape. This suggests a shift in charge accumulation behavior due to enhanced volume change in the electrolyte, allowing additional ionic accommodation and thus boosting capacitance at high fields. Physically, these trends arise because compressibility directly influences the local number density and, hence, the space charge distribution. Unlike the incompressible case, where the number density is constant, compressible systems allow local volume expansion, enabling more ions to be packed near the interface under strong fields. This flexibility affects both the total charge stored and the rate at which it changes with applied voltage.
@@ 5.3 Many constituent mixtures
where $N\neq 3$. This expansion necessitates solving $N-1$ partial mass balance equations in addition to the equations for electric potential and pressure, resulting in a total of $N+1$ coupled partial differential equations.
@@ 6 Concluding remarks and outlook
In this study, we presented a finite element solver for a thermodynamically consistent electrolyte model that rigorously captures the complex interplay between ionic transport, electrostatic interactions, and pressure dynamics in multicomponent electrolyte mixtures. The model is grounded in the principles of non-equilibrium thermodynamics and ensures strict adherence to mass conservation, charge neutrality, and entropy production. Unlike classical models such as the Nernst–Planck system, the proposed framework incorporates steric effects, solvation, and pressure coupling in a physically meaningful and mathematically consistent manner. Despite the model’s increased complexity, the numerical framework exhibits excellent convergence and stability properties, supported by a modular variational formulation that is well-suited for further extensions.
@@
The aim of present study to develop a finite element solver that provides a versatile and physically consistent tool for simulating multicomponent electrolyte systems across diverse applications such as energy storage, water purification, biological transport, and microfluidics. Future work will focus on extending the model to include fluid flow for electrokinetic effects, temperature gradients for thermos-diffusion, and electrochemical reactions at interfaces. Enhancing computational performance through adaptive meshing and integrating the solver into multi-physics frameworks will further broaden its applicability to complex, coupled electrochemical systems.
@@ Appendix A Appendix: Nomenclature
$e_{0}$ Elementary charge [$1.602\times 10^{-19}$ C]
$\varepsilon_{0}$ Vacuum permittivity [$8.85\times 10^{-12}$ F/m]
$\mu_{0}$ Magnetic permeability of vacuum [$4\pi\times 10^{-7}$ N/A2]
$k$ Boltzmann constant [$1.381\times 10^{-23}$ J/K]
$N_{A}$ Avogadro constant [$6.022\times 10^{23}$ 1/mol]

### v2

@@ 3.2 Variational formulation and boundary implementation
$V^{(w)}$ $=\left\{u\in H^{1}(\Omega)\ \middle|\ u=g_{w}\text{ on }\Gamma_{D}^{(w)}\right\},\quad\quad V_{0}={\left\{u\in H^{1}(\Omega)\ \middle|\ u=0\text{ on }\partial\Omega\right\}},$ (39)
@@ 4 Convergence study and validation
[…]
@@ 4.1 Convergence study
Figure 2 illustrates the error measured in the $e_{2}$- and $e_{\infty}$-norms as a function of the number of elements ($n_{x}$) on a logarithmic scale. Both norms exhibit second-order convergence, which is the behavior expected when using first-order elements. Although the FEniCS framework supports higher-order discretization schemes (and extension to such elements is straightforward), we restrict our analysis to the first-order elements specified because they provide sufficient accuracy for the present study. The convergence of both error norms toward zero confirms the solver’s consistency and its ability to approximate the reference solution across the entire domain with increasing resolution.
@@ 5.1.1 Comparison to the classical NP model
The thermodynamically consistent electrolyte model utilized in this paper (hereafter referred to as the DGM model), following [10], addresses key limitations of the classical Nernst–Planck (NP) model and its Poisson–Boltzmann formulation. One central distinction is the proper coupling of diffusion fluxes. In the DGM model, only N1N-1 fluxes are treated as independent, while the flux of the remaining species is obtained from the constraint
@@
which ensures mass conservation and consistency with the barycentric frame. Consequently, the model formulation leads to N1N-1 coupled diffusion equations instead of $N$ uncoupled ones, as seen in the classical NP framework.
@@
Figure 4 compares the anion atomic fraction profiles computed using the classical NP and thermodynamically consistent DGM models under varying applied potential differences. It clearly illustrates that the NP model yields unrealistic concentrations that violate physical bounds, especially near interfaces. In contrast, the DGM model maintains the solution within admissible limits, preventing divergence and ensuring consistency with thermodynamic constraints. The reason for this behavior lies in the DGM model’s incorporation of the constraint $\sum_{\alpha=1}^{N}y_{\alpha}=1$ (see also Section 2.4) and the logarithmic dependence of the chemical potential on the atomic fractions. As the applied potential increases, the DGM model enforces saturation: the anion fraction $y_{A}$ approaches unity near the electrode, while the other fractions decrease toward zero, but none can exceed the physical bounds $[0,1]$. This is a direct consequence of the entropy term in the chemical potential (see (32)), which becomes singular as any $y_{\alpha}\to 0$ or $y_{\alpha}\to 1$. In contrast, the classical NP model lacks this coupling and constraint, allowing concentrations to diverge exponentially with increasing potential, which is nonphysical.
Additionally, Figure 5 presents a broader comparison of electric potential ($\varphi$), pressure ($p$), and atomic fraction ($y_{\alpha}$) profiles at applied voltages $\delta\varphi=1$ and $\delta\varphi=10$. As the applied voltage increases, the classical NP model produces diverging ion concentrations, especially for anions and the neutral solvent, while the DGM model exhibits saturation behavior, limiting atomic fractions to the physical range. These discrepancies become more pronounced under strong electric fields, where the DGM model better reflects the physical blocking of ionic species.
@@
Additionally, we note that while the ideal-mixing entropy captures steric saturation at high concentrations and strong fields, it does not account for explicit ionic diameters or nonlocal excluded-volume correlations. Consequently, the model does not capture packing-induced layering (oscillations) in the EDL structure, resulting in non-monotonic profiles in Figs. 46 (and subsequent figures). To recover such effects, one may employ more sophisticated theories using more advanced nonlocal free energy models (see the discussion in Sec. 6).
@@ 5.2.3 Effect of bulk modulus
Figure 12 illustrates the effect of the bulk modulus KK—which controls the compressibility of the electrolyte—on the one-dimensional electrostatic charge and differential double layer capacitance. The simulations are carried out for increasing compressibility, corresponding to decreasing $K$ values: $K=5\times 10^{4},4\times 10^{4},3\times 10^{4}$, $2\times 10^{4}$, with the incompressible limit ($K\to\infty$). All cases assume a fixed low molarity of $M=0.01\,\text{mol}/\text{L}$. Fig. 12(a) shows that the total electrostatic charge $Q$ increases monotonically with the applied potential difference $\delta\varphi$ for all values of $K$. As expected, compressibility affects the system’s ability to accumulate space charge: lower values of $K$ (i.e., higher compressibility) lead to slightly higher stored charge at the same potential. The differences are small near $\delta\varphi=0$, but they grow with increasing magnitude of the applied potential. As $K$ increases, the results converge smoothly toward the incompressible solution, confirming the theoretical expectation. The more pronounced effect of compressibility appears in the capacitance behavior, shown in Fig. 12(b). For the incompressible case and high $K$ values, the capacitance exhibits a distinct camel-shaped profile with two symmetric peaks and a local minimum at $\delta\varphi=0$. This shape reflects the steric saturation of ions near the electrodes at high field strengths. However, as compressibility increases (i.e., with lower $K$), the profile evolves: the twin humps become broader and less pronounced, and for sufficiently low $K$ (e.g., $K=2\times 10^{4}$), the capacitance begins to increase again beyond the classical humps, deviating from the camel shape. This suggests a shift in charge accumulation behavior due to enhanced volume change in the electrolyte, allowing additional ionic accommodation and thus boosting capacitance at high fields. Physically, these trends arise because compressibility directly influences the local number density and, hence, the space charge distribution. Unlike the incompressible case, where the number density is constant, compressible systems allow local volume expansion, enabling more ions to be packed near the interface under strong fields. This flexibility affects both the total charge stored and the rate at which it changes with applied voltage.
@@ 5.3 Many constituent mixtures
where $N\neq 3$. This expansion necessitates solving N1N-1 partial mass balance equations in addition to the equations for electric potential and pressure, resulting in a total of $N+1$ coupled partial differential equations.
@@ 6 Concluding remarks and outlook
In this study, we presented a finite element solver for a thermodynamically consistent electrolyte model that rigorously captures the complex interplay between ionic transport, electrostatic interactions, and pressure dynamics in multicomponent electrolyte mixtures. The model is grounded in the principles of non-equilibrium thermodynamics and ensures strict adherence to mass conservation, charge neutrality, and entropy production. Unlike classical models such as the Nernst–Planck system, the framework captures steric saturation (via ideal mixing and volume-fraction bounds), solvation-induced volume changes, and pressure coupling; explicit hard-core diameters and nonlocal packing correlations are not included. Despite the model’s increased complexity, the numerical framework exhibits excellent convergence and stability properties, supported by a modular variational formulation that is well-suited for further extensions.
@@
The aim of present study to develop a finite element solver that provides a versatile and physically consistent tool for simulating multicomponent electrolyte systems across diverse applications such as energy storage, water purification, biological transport, and microfluidics. Future work will focus on extending the model to include fluid flow for electrokinetic effects, temperature gradients for thermos-diffusion, and electrochemical reactions at interfaces. Further improvements include an explicit treatment of ion sizes and the incorporation of nonlocal correlations. To recover such effects, one may employ more sophisticated theories, e.g., classical DFT with hard-sphere mixtures and weighted-density approximations [42], or integral-equation approaches for finite-size ions, such as HNC/MSA [39], which can yield oscillatory layering and even charge inversion [18]; see also modified Poisson–Boltzmann formulations with finite-size corrections [46]. Enhancing computational performance through adaptive meshing and integrating the solver into multi-physics frameworks will further broaden its applicability to complex, coupled electrochemical systems.
@@ Appendix A Appendix: Nomenclature
$e_{0}$ Elementary charge $1.602\times 10^{-19}$ C]
$\varepsilon_{0}$ Vacuum permittivity $8.85\times 10^{-12}$ F/m]
$\mu_{0}$ Magnetic permeability of vacuum $4\pi\times 10^{-7}$ N/A2]
$k$ Boltzmann constant $1.381\times 10^{-23}$ J/K]
$N_{A}$ Avogadro constant $6.022\times 10^{23}$ 1/mol]